OpenStudy (tal_tal):

This is a kinda advanced question and I have 3 more but I will fan and medal! Please help! (02.08A)A hyena can run one over four mile in 22.5 seconds. Which of the following correctly shows this rate as miles per hour? one-fourth mile over twenty-two point five seconds = 40 miles per hour twenty-two point five seconds over one-fourth mile = 40 miles per hour one-fourth mile over twenty-two point five seconds = 90 miles per hour twenty-two point five seconds over one-fourth mile = 90 miles per hour Thank you!

OpenStudy (tal_tal):

@cliffordtapp

OpenStudy (mrnood):

you can work put units into an equation in the same way as numbers so look at this as a similar example: I can run 100m in 15 seconds -how fast is that in miles per hour? \[\frac{ 100m }{ 15s } = \frac{ x miles }{ hour }\] 1mile = 1609 m 1 hour = 3600 seconds so \[\frac{ 100m }{ 15s} \times \frac{ 3600s }{1 \min } \times \frac{ 1 mile }{ 1609m }\] you can see that the units cancle out to give miles/hour so multiply out th efractions to get your answer you just need to do the same with YOUR numbers

OpenStudy (mrnood):

! CORRECTION! middle term should be 3600s /1 HOUR

OpenStudy (tal_tal):

so I would put 1/4 a mile over 22.5 seconds?

OpenStudy (tal_tal):

I still don't really get it but oh well lol

OpenStudy (mrnood):

yes 1/4 mile/22.5s (that is the same as 1mile/4*22.5s) then put the conversion from seconds to hours (1hour = 3600s)

OpenStudy (tal_tal):

Wait hold on I think I get it. You put the mile on top and the seconds on the bottom

OpenStudy (tal_tal):

oooh now I get it! Thank you very much!!!

OpenStudy (mrnood):

np - well done